# GeoBEV: Learning Geometric BEV Representation for Multi-view 3D Object Detection

### v1

@@ 3.3 In-Box Label
where $N$ is the total number of GT boxes. It means $p$ is regarded as positive if it is within any GT boxes. From Fig. 4(b), it can be found that the In-Box Label describes the actual geometric structure of objects well and can lead the network to fill the GT boxes with valid features in BEV space as shown in Fig. 1(e).
Although the vanilla In-Box Label reflects the actual geometric structure of objects well, it may cause mismatches between image features and BEV representation of objects. For instance, since Object A in Fig. 4(c) is closer to the camera than Object B, the image records the information of Object A when occlusion occurred (occlusion region is shown by red dotted box). If the blue crosses are regarded as positive, the network will be encouraged to give a high depth score there and mix the information of Object A and Object B, which is harmful for precise detection. A better way is not to supervise the pseudo-points within the occluded region and let the network learn to give a proper depth score by itself. A similar situation happens when objects have irregular shapes, as shown in Fig. 4(d). Not all pseudo-points within the box record the information object and they should also be ignored during training. We use the HTC (Chen et al. 2019) pre-trained on nuImages (Caesar et al. 2020) to provide the mask of objects and filter out the background pseudo-points within GT boxes.
As for the background regions where no GT boxes are available, the LiDAR label is still utilized to supervise the depth scores. It is to make the network learn the whole depth distribution of the scene and locate objects more precisely. Since the LiDAR label reflects the depth of the surfaces while In-Box Label records the actual spatial distribution, we modify the LiDAR label to resolve this divergence between the optimization direction of foreground and background. As shown in Fig. 4(e), we also ignore the pseudo-points behind the background surface, which are used to be regarded as negative. It lets the network adaptively predict how “thick” is the ground and the surrounding buildings and also alleviates the imbalance between the scale of positive and negative.
Method Backbone Image Size Frames mAP NDS mATE mASE mAOE mAVE mAAE
@@
BEVDet4D (Huang and Huang 2022a) ResNet50 256$\times$704 2 0.322 0.457 0.703 0.278 0.495 0.354 0.206
PETRv2 (Liu et al. 2023a) ResNet50 256$\times$704 2 0.349 0.456 0.700 0.275 0.580 0.437 0.187
@@
DETR3D (Wang et al. 2022b) ResNet101 900$\times$1600 1 0.349 0.434 0.716 0.268 0.379 0.842 0.200
PETR (Liu et al. 2022) ResNet101 512$\times$1408 1 0.366 0.441 0.717 0.267 0.412 0.834 0.190
PETRv2 (Liu et al. 2023a) ResNet101 900$\times$ 1600 2 0.421 0.524 0.681 0.267 0.357 0.377 0.186
@@
CAPE (Xiong et al. 2023) ResNet101 512$\times$1408 2 0.431 0.533 - - - - -
STS (Wang et al. 2022c) ResNet101 512$\times$1408 2 0.431 0.542 0.525 0.262 0.380 0.369 0.204
SA-BEV (Zhang et al. 2023a) ResNet101 512$\times$1408 2 0.441 0.549 - - - - -
@@
Table 1: Comparison with previous state-of-the-art multi-view 3D detectors on the nuScenes val set. means the interval between frames may be longer than 0.5s.
3.4 Centroid-Aware Inner Loss
@@
When replacing the LiDAR label with the In-Box Label, the former depth loss should also be changed. We propose Centroid-Aware Inner Loss (CAI Loss) to cooperate with the characteristics of the In-Box Label, which also encourages the model to learn the inner structure of the objects.
First of all, the activation function of the depth scores is changed. The former methods use softmax as the activation function to match the one-hot LiDAR label. It centralizes the depth score on one specific value by:
$\hat{d}_{i}=\frac{e^{d_{i}}}{\sum_{j=1}^{D}e^{d_{j}}},$ (4)
where $d$ is the raw depth score and $\hat{d}$ is the activated one. When the In-Box Label is utilized, the network should give all the pseudo-points within the GT boxes high depth scores. We choose sigmoid as the activation function to independently normalize the depth scores of each pseudo-point within $[0,1]$ by:
$\hat{d}_{i}=\frac{1}{1+e^{-d_{j}}}.$ (5)
The type of depth loss is also changed. When using the LiDAR label to supervise the depth scores, the discrete depth values are regarded as different classes and the depth loss is the mean cross entropy loss of every element of image features. But when using the In-Box Label, it turns into a binary classification of every pseudo-points and severe category imbalance occurs. As a result, focal loss (Lin et al. 2017) is adopted to balance the loss of the different classes to achieve better accuracy. The loss can be calculated by:
$\mathcal{L}_{focal}(p,y)=\begin{cases}-(1-\alpha)p^{\gamma}\log(1-p),&y=0\\ -\alpha(1-p)^{\gamma}\log(p),&y=1\end{cases},$ (6)
where $y$ and $p$ represent the label and the predicted score, $\alpha$ and $\gamma$ are the adjustable parameters.
@@
Besides the spatial distribution of objects, their inner geometric information can also benefit the detection result. Inspired by Centroid-Aware Sampling (Zhang et al. 2022), we give positive pseudo-points different loss weights according to their relative position in the GT boxes. The Centroid-Aware Inner Weight is calculated by:
$W_{CAI}=\sqrt[3]{\frac{\min(f,b)}{\max(f,b)}\times\frac{\min(l,r)}{\max(l,r)}\times\frac{\min(u,d)}{\max(u,d)}},$ (7)
@@
where $f,b,l,r,u,d$ represent the distance of a pseudo-point to the front, back, left, right, up and down surfaces of the GT box, respectively. Only the weight of positive pseudo-points needs to be calculated and the pseudo-point closer to the centroid of an object will have a higher weight as shown in Fig. 4(f). The weights are directly multiplied over the focal loss of positive pseudo-points and the CAI Loss is calculated by:
$\mathcal{L}_{CAI}(p,y)=\begin{cases}-(1-\alpha)p^{\gamma}\log(1-p),&y=0\\ -W_{CAI}\alpha(1-p)^{\gamma}\log(p),&y=1\end{cases}.$ (8)
@@
Using the weighted loss will encourage the pseudo-points near the object centroids to have higher depth scores than the ones near the GT box surfaces. The inner geometric information is thus represented by the depth score differences.
4 Experiments
4.1 Dataset and Metrics
@@
We evaluate our proposed method on the nuScenes (Caesar et al. 2020) dataset, a commonly used large-scale autonomous driving benchmark. It contains diverse 1000 scenarios collected from the real world, each lasting for around 20 seconds. The key samples are annotated at 2Hz and each sample is provided with the data collected from six cameras around the car, one LiDAR on the top of the car and five radars. The 1000 scenarios are split into training set (750 scenarios), validation (150 scenarios) and test set (150 scenarios). The main metric of the nuScenes dataset for 3D object detection is the nuScenes Detection Score (NDS). Except for the commonly used mean average precision (mAP), NDS is also related to five metrics that only take true positive objects into account, including mean Average Translation Error (mATE), mean Average Scale Error (mASE), mean Average Orientation Error (mAOE), mean Average Velocity Error (mAVE), mean Average Attribute Error (mAAE).
Method Backbone Image Size Frames mAP NDS mATE mASE mAOE mAVE mAAE
FCOS3D (Wang et al. 2021) ResNet101 900$\times$1600 1 0.358 0.428 0.690 0.249 0.452 1.434 0.124
DETR3D (Wang et al. 2022b) VoVNet-99 900$\times$1600 1 0.412 0.479 0.641 0.255 0.394 0.845 0.133
@@
PETRv2 (Liu et al. 2023a) VoVNet-99 640$\times$1600 2 0.490 0.582 0.561 0.243 0.361 0.343 0.120
@@
Table 2: Comparison with previous state-of-the-art multi-view 3D detectors on the nuScenes test set. means the interval between frames may be longer than 0.5s.
Method Backbone Image Size Frames mAP NDS mATE mASE mAOE mAVE mAAE
BEVPoolv2 (Huang and Huang 2022b) ResNet50 256$\times$704 9 0.406 0.526 0.572 0.275 0.463 0.275 0.188
VideoBEV (Han et al. 2023) ResNet50 256$\times$704 8 0.422 0.535 0.564 0.276 0.440 0.286 0.198
Sparse4Dv2 (Lin et al. 2023) ResNet50 256$\times$704 - 0.439 0.539 0.598 0.270 0.475 0.282 0.179
Table 3: Comparison with the long-term temporal methods on the nuScenes val set. Benefited from the perspective-view pre-training.
4.2 Implementation Details
We adopt the BEVDepth (Li et al. 2023c) as the baseline and add our proposed In-Box Label and Radial-Cartesian BEV Sampling to build GeoBEV. When comparing with state-of-the-art methods, we follow the commonly used configurations. For the experiments on the nuScenes validation set, the ResNet50 and ResNet101 are adopted as the backbone to process the multi-view image downsampled to 256$\times$704 and 512$\times$1408, respectively. When evaluating on the nuScenes test set, the VoVNet-99 (Lee et al. 2019) pre-trained by DD3D (Park et al. 2021) is adopted as the backbone to process the images cropped to 640$\times$1600. These models are trained for 20 epochs with CBGS strategy (Zhu et al. 2019). Except for regular data augmentation in the image and BEV space, the BEV-Paste strategy (Zhang et al. 2023a) is also used to alleviate overfitting during the long training process. For the ablation study, we use ResNet50 (He et al. 2016) as the image backbone and the image size is downsampled to 256$\times$704. The models are trained for 24 epochs without the CBGS strategy.
4.3 Main Results
We compare GeoBEV with previous state-of-the-art multi-view 3D detectors on the nuScenes val and test set. The experiment results in Tab. 1 show that GeoBEV achieves the best detection accuracy on nuScenes val set. When detecting from images in 256$\times$704 and using ResNet50 as the backbone, GeoBEV outperforms SA-BEV (Zhang et al. 2023a), the previous state-of-the-art, by a significant margin of 2.8% mAP and 2.3% NDS. When increasing image resolution to 512$\times$1408 and using ResNet101 as the backbone, GeoBEV stays ahead of the curve and outperforms SA-BEV by 3.8% mAP and 3.3% NDS.
The results of experiments made on the nuScenes test set are shown in Tab. 2. It can be found that GeoBEV also gets the best performance and outperforms SA-BEV by 1.0% mAP and 1.1% NDS. Those persuasive experiment results highlight the effectiveness of GeoBEV.
We also conduct experiments to figure out whether the BEV representations with high-quality geometric information can combine with the long-term temporal fusion strategy and show the results in Table 3. When using 7 past frames with an interval of 0.5 second to supplement the current scenes, GeoBEV exceeds state-of-the-art long-term temporal methods and outperforms StreamPETR (Wang et al. 2023) by 2.9% mAP and 2.5% NDS.
4.4 Ablation Study
@@
We evaluate the contributions of our proposed components and show the results in Tab. 4. It can be found that both RC-Sampling and In-Box Label effectively increase the detection accuracy. RC-Sampling increases 2.6% mAP and 3.3% NDS compared to the baseline. Applying In-Box Label and CAI Loss also boosts the performance by 2.2% mAP and 2.2% NDS. After combining the two components, the performance is increased by 4.4% mAP and 4.4% NDS in total.
@@
To evaluate the capacity of RC-Sampling, we compare it with the most efficient LSS-based and Transformer-based feature transformation methods. BEVPoolv2 (Huang and Huang 2022b) and DFA3D (Li et al. 2023a) are chosen as the representatives respectively. From the experiment results in Tab. 5, it can be found the detection accuracy of RC-Sampling outperforms both BEVPoolv2 and DFA3D in generating BEV representations with different resolutions. Besides, RC-Sampling exhibits better real-time performance and achieves the best FPS under the same configuration as other methods. The resolution of depth scores also influences the geometric quality of BEV representations, so we upsample the size of depth scores by utilizing several convolutions to provide fine-grained geometry information to RC-Sampling, which further increases the performance by 0.5% mAP and 0.7% NDS without significantly affects its efficiency.
@@
We conduct experiments to evaluate different configurations when applying the In-Box Label as in Tab. 6. When simply replacing the LiDAR label with the vanilla In-Box Label, the performance is increased by 0.8% mAP and 0.8% NDS. It is further improved by 0.2% mAP / 0.2% NDS and 0.4% mAP / 0.4% NDS after using sigmoid as the activation function and letting the depth scores supervised by focal loss. We also compare the performance between vanilla In-Box Label and complete In-Box Label, the results show that the In-Box Label is more in line with the real world and has an advantage of 0.5% mAP and 0.4% NDS. When replacing Focal Loss with Centroid-Aware Inner Loss, there is another 0.3% mAP and 0.4% NDS improvement, which illustrates that inner geometric structure is helpful for the detection.
5 Conclusion
@@
In this paper, we propose a novel multi-view 3D object detector, namely GeoBEV, which generates BEV representation that restores authentic geometric information of the scene. The Radial-Cartesian BEV Sampling simply does high-dimensional matrix multiplication between transposed image features and depth scores to obtain Radial BEV features, which are then transformed into Cartesian BEV features by bilinear sampling. This approach can rapidly generate high-resolution BEV representations while effectively avoiding the presence of vacant feature values. Based on the physics of the real world, In-Box Label can reflect the actual geometric structure of objects, effectively improving the accuracy of the information carried by BEV representation. Centroid-Aware Inner Loss cooperates with In-Box Label to make full of its advantage and also encourages the network to learn the inner geometry of objects.
@@
We conduct extensive experiments on nuScenes dataset and GeoBEV reaches a new state-of-the-art, highlighting the effectiveness of Radial-Cartesian BEV Sampling and In-Box Label. Theoretically, these components can be easily integrated into many existing BEV-based detectors, improving the geometry quality of BEV representation and increasing the detection performance.

### v2

@@ Introduction
where $N$ is the total number of GT boxes. It means if $p$ is within any GT boxes, it is regarded as positive. Such depth labels encourage the model to describe the actual geometric structure of objects well and fill the GT boxes with valid features in BEV space as shown in Fig. 1(d).
However, Vanilla In-Box Label may cause mismatches between image features and BEV representation of objects and several corrections are needed. For instance, since Object A in Fig. 4(a) has a smaller depth than Object B, the image records the information of Object A in the occluded area (represented by the red dotted box). If the blue crosses are treated as positive, the network will give a high depth score there and mix Object B with A, which is harmful to perception. We choose not to supervise the pseudo-points within the occluded region and let the network learn to give a proper depth score by itself. A similar solution is adopted when objects have irregular shapes, as shown in Fig. 4(b). Not all pseudo-points within the GT box record the information object and they should also be ignored during training. We use the HTC (Chen et al. 2019) pre-trained on nuImages (Caesar et al. 2020) to provide the mask of objects and filter out the background pseudo-points while calculating loss.
As for the background regions where no GT boxes are available, the LiDAR label is still employed to make the network learn the whole depth distribution of the scene and locate objects more precisely. Since the LiDAR label reflects the depth of the surfaces while In-Box Label records the actual spatial distribution, we modify the LiDAR label to resolve the optimization divergence between foreground and background. As shown in Fig. 4(c), we also ignore the pseudo-points behind the background surface (represented by blue dotted boxes), which are used to be negative. It lets the network adaptively predict how “thick” is the ground and the surrounding buildings and also balances the number of positive and negative.
[…]
@@
PETRv2 (Liu et al. 2023) ResNet50 256$\times$704 2 0.349 0.456 0.700 0.275 0.580 0.437 0.187
@@
BEVNext (Li et al. 2024) ResNet50 256$\times$704 8 0.456 0.560 0.530 0.264 0.424 0.252 0.206
RayDN (Liu et al. 2024) ResNet50 256$\times$704 8 0.469 0.563 0.579 0.264 0.433 0.256 0.187
@@
PETRv2 (Liu et al. 2023) ResNet101 900$\times$ 1600 2 0.421 0.524 0.681 0.267 0.357 0.377 0.186
@@
SOLOFusion (Park et al. 2022) ResNet101 512$\times$1408 17 0.483 0.582 0.503 0.264 0.381 0.246 0.207
StreamPETR (Wang et al. 2023) ResNet101 512$\times$1408 8 0.504 0.592 0.569 0.262 0.315 0.257 0.199
BEVNext (Li et al. 2024) ResNet101 512$\times$1408 8 0.500 0.597 0.487 0.260 0.343 0.245 0.197
RayDN (Liu et al. 2024) ResNet101 512$\times$1408 8 0.518 0.604 0.541 0.260 0.315 0.236 0.200
@@
Table 1: Comparison with previous state-of-the-art multi-view 3D detectors on the nuScenes val set. Benefited from the perspective-view pre-training.
Centroid-Aware Inner Loss
@@
Adapted to the characteristics of In-Box Label, we propose Centroid-Aware Inner Loss (CAI Loss) to replace the Cross-Entropy depth loss. It encourages the model to learn the inner structure of the objects and further refine the geometric information of BEV representation.
Softmax and Cross-Entropy Loss are formally chosen as the activation function and the depth loss to match the one-hot LiDAR label. Softmax centralizes the depth score on the depth values of the object surfaces and Cross-Entropy Loss treats discrete depth values as different classes. When In-Box Label is utilized, the network should give all the pseudo-points within the GT boxes high depth scores. We choose Sigmoid to independently normalize the depth scores of each pseudo-point within $[0,1]$. Besides, the multiple classification of discrete depth values turns into the binary classification of whether pseudo-points are in boxes, which results in far more negative than positive. As a result, Focal Loss (Lin et al. 2017) is adopted to balance the losses of different classes.
@@
To learn the inner structure of objects, we vary the loss weights of positive pseudo-points according to their relative position in the GT boxes. Inspired by  (Zhang et al. 2022), Centroid-Aware Inner Weight is defined as:
$W_{\scriptscriptstyle CAI}=\sqrt[3]{\frac{\min(f,b)}{\max(f,b)}\times\frac{\min(l,r)}{\max(l,r)}\times\frac{\min(u,d)}{\max(u,d)}},$ (6)
@@
where $f,b,l,r,u,d$ represent the distance of a pseudo-point to the front, back, left, right, up and down surfaces of the GT box. Only the weight of positive pseudo-points needs to be calculated and the pseudo-point closer to the centroid of an object will have a higher weight as shown in Fig. 4(d). The weights are directly multiplied over the focal loss of positive pseudo-points and the CAI Loss is calculated by:
$\mathcal{L}_{\scriptscriptstyle CAI}(p,y)=\begin{cases}-(1-\alpha)p^{\gamma}\log(1-p),&y=0\\ -W_{\scriptscriptstyle CAI}\alpha(1-p)^{\gamma}\log(p),&y=1\end{cases},$ (7)
@@
where $y$ and $p$ are the label and the activated depth score, $\alpha$ and $\gamma$ are the parameters of Focal Loss. CAI loss will let the pseudo-points near the object centroids have higher depth scores than the ones near the GT box surfaces, thus expressing the inner geometric information of objects.
Experiments
Dataset and Metrics
@@
We evaluate our proposed method on the nuScenes (Caesar et al. 2020) dataset, a commonly used autonomous driving benchmark. It contains 1000 scenarios collected from the real world, each lasting for around 20 seconds. The key samples are annotated at 2Hz and each sample is provided with the data collected from six cameras, one LiDAR and five radars. The 1000 scenarios are split into training set (750 scenarios), validation (150 scenarios) and test set (150 scenarios). The main metric of the nuScenes dataset for 3D object detection is the nuScenes Detection Score (NDS). Except for the commonly used mean average precision (mAP), NDS is also related to five metrics that only take true positive objects into account, including mean Average Translation Error (mATE), mean Average Scale Error (mASE), mean Average Orientation Error (mAOE), mean Average Velocity Error (mAVE), mean Average Attribute Error (mAAE).
[…]
@@
PETRv2 (Liu et al. 2023) VoVNet-99 640$\times$1600 2 0.490 0.582 0.561 0.243 0.361 0.343 0.120
@@
StreamPETR (Wang et al. 2023) VoVNet-99 640$\times$1600 8 0.550 0.636 0.479 0.239 0.317 0.241 0.119
BEVNext (Li et al. 2024) VoVNet-99 640$\times$1600 8 0.557 0.642 0.409 0.241 0.352 0.233 0.129
OPEN (Hou et al. 2025) VoVNet-99 640$\times$1600 - 0.567 0.644 0.456 0.244 0.325 0.240 0.129
RayDN (Liu et al. 2024) VoVNet-99 640$\times$1600 8 0.565 0.645 0.461 0.241 0.322 0.239 0.114
@@
GeoBEV VoVNet-99 640$\times$1600 8 0.579 0.662 0.369 0.234 0.323 0.229 0.120
Table 2: Comparison with previous state-of-the-art multi-view 3D detectors on the nuScenes test set.
Implementation Details
We adopt the BEVDepth (Li et al. 2023c) as the baseline to build GeoBEV and compare it with state-of-the-art methods in the commonly used configurations. For the experiments on the nuScenes validation set, the ResNet50 and ResNet101 (He et al. 2016) are adopted as the backbone to process the images in 256$\times$704 and 512$\times$1408, respectively. When evaluating on the nuScenes test set, the VoVNet-99 (Lee et al. 2019) pre-trained by DD3D (Park et al. 2021) is adopted as the backbone to process the images cropped to 640$\times$1600. These models are trained for 20 epochs with CBGS strategy (Zhu et al. 2019). Except for regular data augmentation, the BEV-Paste (Zhang et al. 2023a) is adopted to alleviate overfitting during the long training process. Future frames and test-time augmentation are not adopted. For the ablation study, we use ResNet50 as the image backbone and the models are trained for 24 epochs without the CBGS strategy.
Main Results
We compare GeoBEV with previous state-of-the-art multi-view 3D detectors on the nuScenes val and test set. The experiment results in Tab. 1 show that GeoBEV achieves the best detection accuracy on nuScenes val set at different configurations. When detecting from images in 256$\times$704 and using ResNet50 as the backbone, GeoBEV outperforms RayDN (Liu et al. 2024), the previous state-of-the-art, by 1.0% mAP and 1.2% NDS. When increasing image resolution to 512$\times$1408 and using ResNet101 as the backbone, GeoBEV stays ahead of the curve and outperforms RayDN by 0.8% mAP and 1.1% NDS.
The experiment results on the nuScenes test set are shown in Tab. 2. GeoBEV also gets the best performance of 57.9% mAP / 66.3% NDS, surpassing StreamPETR (Wang et al. 2023) by 2.9% mAP / 2.6% NDS and RayDN by 1.4% mAP / 1.7% NDS, respectively. Those persuasive experiment results highlight the effectiveness of GeoBEV.
Ablation Study
@@
We evaluate the contributions of our proposed components and show the results in Tab. 3. It can be found that both RC-Sampling and In-Box Label effectively increase the detection accuracy. When using BEVDepth (Li et al. 2023c) as the baseline, there is an improvement of 2.6% mAP and 3.3% NDS after applying RC-Sampling. In-Box Label and CAI Loss also boost the performance by 2.2% mAP and 2.2% NDS. After combining the two components, the performance is increased by 4.4% mAP and 4.4% NDS in total. To estimate the versatility of our proposed components, we also choose BEVDet (Huang et al. 2021) and BEVStereo (Li et al. 2023b) as the baselines. After adopting RC-Sampling and In-Box Label, their accuracy is improved by 2.7% mAP / 4.1% NDS and 3.4% mAP / 3.9% NDS respectively.
@@
To show the capacity of RC-Sampling, we compare it with the most efficient LSS-based and Transformer-based feature transformation methods. BEVPoolv2 (Huang and Huang 2022b) and DFA3D (Li et al. 2023a) are chosen as the representatives. The comparison with Voxel-Sampling, the unoptimized version of RC-Sampling, is also implemented. All of the feature transformation methods are incorporated into the same BEVDepth model. From the experiment results in Tab. 4, it can be found the detection accuracy of RC-Sampling outperforms both BEVPoolv2 and DFA3D, indicating the better geometric quality of BEV representation. Besides, RC-Sampling exhibits better real-time performance and achieves the best FPS while generating BEV representation with different resolutions. Voxel-Sampling achieves comparable accuracy as RC-Sampling, but its speed is far behind. More details of the comparison with Voxel-Sampling are shown in the supplementary material. We also upsample the size of depth scores by lightweight convolution to provide fine-grained geometry information to RC-Sampling, which further increases the performance by 0.5% mAP and 0.7% NDS without significantly affecting its efficiency.
@@
We conduct experiments to evaluate different configurations when applying the In-Box Label as in shown Tab. 5. When simply replacing the LiDAR label with the Vanilla In-Box Label, the performance is increased by 0.8% mAP and 0.8% NDS. It is further improved by 0.2% mAP / 0.2% NDS and 0.4% mAP / 0.4% NDS after using Sigmoid as the activation function and letting the depth scores supervised by Focal Loss. We also compare the performance between Vanilla In-Box Label and the complete In-Box Label, the results show that the In-Box Label is more in line with the real world and has an advantage of 0.5% mAP and 0.4% NDS. When replacing Focal Loss with Centroid-Aware Inner Loss, there is another 0.3% mAP and 0.4% NDS improvement, which illustrates that inner geometric structure is helpful for the detection.
Conclusion
@@
In this paper, we propose a novel multi-view 3D object detector, namely GeoBEV, which generates BEV representation that restores authentic geometric information of the scene. Radial-Cartesian BEV Sampling simply does high-dimensional matrix multiplication between transposed image features and depth scores to obtain Radial BEV features, which are then transformed into Cartesian BEV features by bilinear sampling. This approach can rapidly generate high-resolution BEV representation while effectively avoiding the presence of vacant feature values. Based on the physics of the real world, In-Box Label can reflect the actual geometric structure of objects, effectively improving the accuracy of the information carried by BEV representation. Centroid-Aware Inner Loss cooperates with In-Box Label to make full of its advantage and also encourages the network to learn the inner geometry of objects.
@@
We conduct extensive experiments on nuScenes dataset and GeoBEV reaches a new state-of-the-art, highlighting the effectiveness of our proposed components in enhancing the geometric quality of BEV representation. Additional experiments also illustrate that these components can be easily integrated into many existing BEV-based detectors and bring stable improvement in accuracy and real-time performance.
Acknowledgment
This research was supported by Zhejiang Provincial Natural Science Foundation of China under Grant No. LD24F020016 and LQ23F020024, National Natural Science Foundation of China under Grant No. 62302031, and “Pioneer” and “Leading Goose” R&D Program of Zhejiang under Grant No. 2024C01020.
Appendix A More Implementation Details
We implement experiments of GeoBEV with 8 NVIDIA GeForce RTX 3090 GPUs. AdamW (Loshchilov and Hutter 2017) with a decay weight of 0.01 is adopted as the optimizer and the learning rate is set to $2\times 10^{-4}$. We apply data augmentation in both image space and BEV space. The multi-view images are processed by random scaling with a range of $[0.86,1.25]$ and horizontal flipping with a probability of 0.5. When training with CBGS (Zhu et al. 2019), the range of random scaling is expanded to $[0.5,1.25]$. Following BEVDet (Huang et al. 2021), the BEV features are processed by random scaling with a range of $[0.95,1.05]$, random flipping of the X and Y axes with a probability of 0.5 and random rotating with a range of $[-22.5^{\circ},22.5^{\circ}]$. When processing images larger than or equal to $512\times 1408$, the size of BEV features is increased from $256\times 256$ to $512\times 512$.
Appendix B More Experiment Results
More Results on RC-Sampling
We provide more comparison results between RC-Sampling and the unoptimized Voxel-Sampling on time and memory costs in Tab. 6. Except for BEV size, the height of the voxel features also influences the efficiency of Voxel-Sampling. We partition the space by different numbers of heights and the experiment results show that higher height resolution leads to higher performance but also significantly increases the time and memory costs. On the contrary, RC-Sampling accumulates the features along the height before the sampling operations, which restores more complete height information than Voxel-Sampling by much less computation. From the experiment results, RC-Sampling achieves the same accuracy performance as Voxel-Sampling with a height resolution of 20 while spending less than 10% of the time and memory costs. When producing the larger BEV representation, the time and memory costs of Voxel-Sampling become impracticable, while RC-Sampling remains deployment-friendly. Compared with BEVPoolv2 (Huang and Huang 2022b), RC-Sampling shows advantages in efficiency and accuracy performance while slightly increasing the memory cost.
More Results on In-Box label
Tab. 7 shows more detailed results of the ablation study on the In-Box label. By default, the depth scores in the background region, where GT boxes are not available, are still supervised by the LiDAR label. If all background pseudo-points are regarded as negative classes, the requirement for LiDAR points during the training process will be eliminated, thereby significantly reducing data acquisition costs. Compared to the model that does not use any depth labels, only using In-Box label increases 1% mAP and 1.2% NDS. The results in Tab. 7 also show the effectiveness of several optimizations applied in In-Box label, including ignoring the pseudo-points in occlusion, the background pseudo-points within GT boxes and the pseudo-points behind the background surface. It illustrates that the optimized In-Box label which is more in line with the real world can lead to more precise detection.
[…]
Voxel-Sampling $128\times 128$ 5 0.337 0.458 1.77ms 204.1MB
$128\times 128$ 10 0.343 0.460 2.84ms 252.1MB
$128\times 128$ 20 0.342 0.464 3.18ms 352.1MB
$256\times 256$ 20 0.354 0.484 11.97ms 952.1MB
BEVPoolv2 $128\times 128$ 0.337 0.456 0.43ms 19.5MB
$256\times 256$ 0.344 0.474 0.62ms 48.7MB
RC-Sampling $128\times 128$ 0.344 0.465 0.29ms 20.0MB
$256\times 256$ 0.358 0.482 0.53ms 49.5MB
Table 6: More comparison between different feature transformation modules. “H” denotes the height of the voxel features created by Voxel-Sampling. “Latency” and “Memory” are cost only by feature transformation approaches.
[…]
None 0.330 0.454
In-Box 0.351 0.470
0.340 0.466
0.353 0.472
0.353 0.473
0.356 0.474
Table 7: More ablation study on the In-Box label. “w/o LiDAR” denotes not using LiDAR points during training. “$O_{A}$” means ignoring the pseudo-points in occlusion. “$O_{B}$” means ignoring the background pseudo-points within GT boxes. “$O_{C}$” means ignoring the pseudo-points behind the background surface.
Figure 5: Visualization results on BEV representation of GeoBEV and its baseline. The background is filtered out to show the difference in the foreground. The red boxes and green boxes represent the ground truth and the predicted boxes, respectively. The dashed blue rectangles illustrate that the geometry guided BEV representations result in higher detection accuracy.
Appendix C Visualization
We present the visualization results of GeoBEV and its baseline in Fig. 5. The L1 norm of each position in BEV features is first calculated. After that, brightness positively correlated with the L1 norm is assigned in the figure. To show the difference in the foreground, we employ a pretrained instance segmentation model to provide foreground masks on the images. These masks are used to filter out the background regions before transforming image features into BEV features.
It can be found that the BEV representation of GeoBEV effectively restores the fine-grained geometric structure of the objects. This helps the network more clearly identify the size and orientation of the objects as illustrated by dashed blue rectangles, leading to improved detection performance.